OpenStudy (anonymous):

X - Y=17 ; (8,-9) X + Y=-1

OpenStudy (texaschic101):

x - y = 17 x + y = -1 ----------add 2x = 16 x = 8 x + y = -1 8 + y = -1 y = -1 - 8 y = -9 solution is : (8,-9)